OpenStudy (calculusxy):

MEDAL!!! Need help on finding the vertex of \(y = -3x^2 + 4x\)

OpenStudy (swag889):

Will you give me owl bucks? :D

OpenStudy (calculusxy):

I don't have any ...

OpenStudy (swag889):

=.=

OpenStudy (calculusxy):

I first tried to find the axis of symmetry. \(\large x = \frac{-b}{2a}\) \(\large x = -\frac{4}{-6} \rightarrow -\frac{2}{3}\) I don't know what to do with the fraction. Should I keep it the same or somehow change it?

OpenStudy (swag889):

here http://i.imgur.com/OoIMvCp.png

OpenStudy (swag889):

^ that should help you ^_^

OpenStudy (calculusxy):

So I guess I can keep it?

OpenStudy (calculusxy):

@Nnesha

OpenStudy (superdavesuper):

u can convert to vertex form by "completing the square"; read more from here: http://www.purplemath.com/modules/sqrvertx.htm u actually have got the first step right by finding x-coordinate of the vertex, h. just need to find a and k in the vertex form of y = a(x-h)^2 + k

OpenStudy (calculusxy):

And to find the \(h\) I just need to plug in \(\frac{2}{3}\) to when \(x\) is.

OpenStudy (mathmale):

Towards completing the square: try factoring out -3 from the 2 terms of the given expression; then follow superdavesuper.

OpenStudy (swag889):

Can I get my medal? ._.

OpenStudy (mathmale):

Skip the medal business, would you please?

OpenStudy (swag889):

:O

OpenStudy (superdavesuper):

@calculusxy u found h already but yes, to find k, simply calculate y when x=h

OpenStudy (calculusxy):

Okay.

OpenStudy (mathmale):

@calculusxy: You have found the x-coordinate of the vertex already, using x=-b/(2a). Follow superdave's directions, above.

OpenStudy (superdavesuper):

@calculusxy dont forget what @mathmale said, u need to do that in finding a

OpenStudy (calculusxy):

Okay. I think I might have phrased the question wrong. I need to find the vertex for graphing.

OpenStudy (mathmale):

You made an effort to find the x-coord. of the vertex using x=-b/(2a). You got that right EXCEPT for the sign; the sign is (+).

OpenStudy (calculusxy):

Oh yes! You're right @mathmale

OpenStudy (mathmale):

What is a? What is b? Answer, then calculate x=-b/(2a)

OpenStudy (superdavesuper):

if only the vertex location is needed but not the vertex form, then u are almost done @calculusxy :)

OpenStudy (calculusxy):

\[\large y = -3(\frac{ 2 }{ 3 })^2 + 4(\frac{ 2 }{ 3 }) = -\frac{ 4 }{ 3 } \times \frac{ 8 }{ 3 } = -\frac{ 32 }{ 9 } = -3\frac{ 5 }{ 9 }\]

OpenStudy (mathmale):

I'm not going to check that in every detail, but you're off to a good start.

OpenStudy (calculusxy):

So my vertex would be \(\large (\frac{2}{3}, -3\frac{5}{9})\)

OpenStudy (mathmale):

Now write the coordinates of the vertex. You've just done that. Looks good. Proper format and all that. Nice work. Thanks for y our patience. Over and out.

OpenStudy (mathmale):

I'm off the 'Net for the time being.

OpenStudy (calculusxy):

Thank you @superdavesuper and @mathmale !!!

OpenStudy (superdavesuper):

@calculusxy u may want to double-check....the wolf said h is wrong... http://www.wolframalpha.com/input/?i=y+%3D+-3x%5E2+%2B+4x

OpenStudy (calculusxy):

Where do you see that?

OpenStudy (calculusxy):

@superdavesuper

OpenStudy (superdavesuper):

clock on the property button ;)

OpenStudy (superdavesuper):

click i mean

OpenStudy (calculusxy):

I still don't see it (sorry)

OpenStudy (superdavesuper):

in the table, second row...Geometric figures parabola....scroll to the right - white button labeled properties

OpenStudy (mathmale):

I have double checked and still assert that x=2/3 is the x-coordinate of the vertex. The y-coordinate is simply the value of the given function at x=2/3: y = -3x^2 + 4x becomes k = -3(2/3)^2 + 4(2/3) = ?? Vertex of parabola is (2/3, ?? )

OpenStudy (superdavesuper):

@mathmale yes, x=2/3 is right but y is not at \[-3\frac{ 5}{ 9 }\]

OpenStudy (mathmale):

Agreed. See my previous comment.